# Efficient Latent Variable Causal Discovery: Combining Score Search and Targeted Testing

### v2

@@ Abstract
We present a family of score-guided mixed-strategy causal search algorithms that extend this framework. First, we introduce BOSS-FCI and GRaSP-FCI, variants of GFCI (Greedy Fast Causal Inference) that substitute BOSS (Best Order Score Search) or GRaSP (Greedy Relaxations of Sparsest Permutation) for FGES (Fast Greedy Equivalence Search), preserving correctness while trading off scalability and conservativeness. Second, we develop FCI Targeted-Testing (FCIT), a novel hybrid method that replaces exhaustive testing with targeted, score-informed tests guided by BOSS. FCIT guarantees well-formed PAGs and achieves higher precision and efficiency across sample sizes. Finally, we propose a lightweight heuristic, LV-Lite (Latent Variable “Light”), which returns the PAG of the BOSS DAG (Directed Acyclic Graph). Though not strictly sound for latent confounding, LV-Lite often matches FCIT’s accuracy while running substantially faster.
Simulations and real-data analyses show that BOSS-FCI and GRaSP-FCI provide robust baselines, FCIT yields the best balance of precision and reliability, and LV-Lite offers a fast, near-equivalent alternative. Together, these methods demonstrate that targeted and score-guided strategies can dramatically improve the efficiency and correctness of latent-variable causal discovery.
@@ 1 Introduction
Learning causal structure from observational data is central in fields such as epidemiology, economics, and biology, where randomized experiments are often infeasible. In many real-world settings, latent variables or selection bias complicate this task, undermining the reliability of traditional discovery methods. The Fast Causal Inference (FCI) algorithm [14] provides a principled framework for such cases by applying conditional independence (CI) tests to subsets of adjacent variables (or Possible-D-SEP sets). While correct in theory, this strategy is fragile in practice: testing many subsets amplifies stochastic variation in $p$-values, leading to spurious independence claims, missing or extra edges, and unstable orientations. We refer to this as the repeated testing problem.
@@
Beyond sampling variability, another challenge is almost unfaithfulness [18, 15]. Even when the Causal Markov and Faithfulness assumptions hold exactly, small parameter perturbations can produce distributions that are nearly unfaithful—where partial correlations or other dependence measures are close to zero but not exactly zero. In finite samples, such “almost unfaithful” parameterizations lead to unstable independence decisions: small statistical fluctuations can flip dependence judgments, causing orientation errors or spurious adjacencies. As shown by Zhang and Spirtes, [18], this phenomenon can make reliable causal inference difficult even for consistent algorithms, motivating procedures that are robust to weak dependencies or employ stronger assumptions such as $k$-triangle faithfulness [15]. We therefore treat almost unfaithfulness as a complementary concern to the repeated-testing problem—structural rather than statistical in origin, but equally capable of producing unstable PAGs in practice.
@@
Several extensions have sought to mitigate such issues. Order-independent FCI variants [5] stabilize orientations, and score-based methods [^1] Scores in Tetrad currently cover linear Gaussian and linear non-Gaussian models (using BIC-type criteria), multinomial models, mixed multinomial/linear Gaussian models, and, more recently, mixed multinomial/nonlinear post-nonlinear models via truncated basis-function scores [11]. extend causal search to latent-variable models through likelihood-based criteria [16, 4]. Hybrid approaches such as Greedy FCI (GFCI; [10]) combine score-based search with CI testing, but still rely on adjacency-based phases—i.e., during the adjacency-removal stage, testing all subsets of nodes adjacent to a pair—and thus remain vulnerable to repeated testing and almost unfaithfulness. Constraint-based methods, by contrast, have the advantage of being inherently nonparametric and making no distributional assumptions except for those of the selected conditional independence test, which motivates the option of keeping such testing in the algorithm.
@@
Recent advances in score-based search motivate new hybrids that combine the distributional flexibility of constraint-based methods with the efficiency of likelihood-based optimization. Best Order Score Search (BOSS; [2]) provides a scalable, permutation-based alternative to FGES, while GRaSP [9] offers a related variant. Substituting these into the GFCI framework yields BOSS-FCI and GRaSP-FCI—hybrid algorithms that preserve the nonparametric generality of constraint-based search while leveraging score-based initialization to improve stability and scalability. The original GFCI algorithm [10] is sound and complete in the population limit under the Causal Markov and Faithfulness assumptions, returning the correct Partial Ancestral Graph when provided with perfect independence information. However, in finite samples, GFCI and its variants still depend on adjacency-based conditioning and thus remain susceptible to repeated testing and almost unfaithfulness.
@@
Building on these baselines, we propose FCI Targeted-Testing (FCIT), a mixed-strategy algorithm that replaces exhaustive adjacency-based testing with a path-based procedure. Like GFCI, FCIT begins with a score-based completed partially directed acyclic graph (CPDAG)—a representation of the Markov equivalence class of DAGs that share the same adjacencies and unshielded colliders. Score-based algorithms such as BOSS and GRaSP are score-consistent: in the large-sample limit, they recover the correct CPDAG under the Causal Markov and Faithfulness assumptions for the model class (e.g., linear Gaussian, multinomial, or post-nonlinear) used to define the score. FCIT then applies a recursive routine, block_paths_recursively, to identify conditioning sets only when necessary to block active paths between variables, thereby avoiding the over-conditioning that arises in FCI and GFCI from testing all subsets of adjacent variables. Compared with these algorithms, FCIT performs far fewer conditional independence tests, which reduces sampling error and increases the reliability of the overall learned structure. Together with a well-formedness check that enforces the rules of a Partial Ancestral Graph (PAG)—a mixed graph representing the Markov equivalence class of maximal ancestral graphs consistent with possible latent variables and selection bias—this targeted strategy guarantees that every intermediate graph remains structurally consistent, meaning it obeys the ancestral, acyclicity, and maximality constraints required of a valid PAG. The result is greater efficiency, more stable orientations, and structural correctness not ensured by prior FCI variants [17, 8].
@@
Finally, we introduce a deliberately simple heuristic, LV-Lite (Latent Variables “Light”), which outputs the PAG implied by the BOSS DAG without performing any latent-specific conditional independence testing. Though it cannot orient bidirected edges, LV-Lite scales exceptionally well and, in extensive simulation studies, often matches or exceeds the empirical accuracy of more complex algorithms when compared to the true PAG.
@@
Together, these four algorithms—BOSS-FCI, GRaSP-FCI, FCIT, and LV-Lite—form a continuum. The first two establish sound baselines by embedding advanced score-based methods within the GFCI framework. FCIT introduces a principled path-based refinement that improves efficiency and reliability, while LV-Lite represents a heuristic extreme: formally incomplete but practically effective. Our empirical evaluation, spanning both simulation and real data, demonstrates the complementary strengths of these approaches. BOSS-FCI and GRaSP-FCI retain theoretical correctness but are constrained by repeated testing; FCIT balances correctness with improved precision and scalability; and LV-Lite sacrifices correctness guarantees yet achieves strong accuracy with minimal computational cost in simulation.
@@
The remainder of this paper is organized as follows. Section 2 reviews graphical preliminaries. Sections 34 present the four algorithms. Section 5 introduces optimization details, Section 6 outlines theoretical guarantees, and Section 8 reports empirical results. All algorithms are implemented in the open-source Tetrad suite [12], with reproducibility materials available on GitHub.
@@
  • A heuristic, LV-Lite, that outputs the PAG of the BOSS DAG, achieving high empirical accuracy at minimal cost.
@@ 2 Graphical Preliminaries
In this section, we define key terms related to graphical causal models used throughout the paper. We focus on models based on causal directed acyclic graphs (DAGs), which consist of directed edges $X\rightarrow Y$. Such an edge indicates that variable $X$ causes variable $Y$, meaning that interventions on $X$ counterfactually alter $Y$ [14].
@@
A MAG or PAG is maximal (following Zhang [17]). if, for every non-adjacent pair $X,Y$, there exists a conditioning set S such that $X$ and $Y$ are m-separated given S. Equivalently, there are no inducing paths between non-adjacent nodes.
@@
A set of variables $\textbf{(}V)$ is causally sufficient if it includes all common causes of observed variables in $\textbf{(}V)$; otherwise, it is causally insufficient. A model is edge-minimal of a certain type (e.g., DAG, CPDAG, PAG) if no model of that type has fewer edges while remaining Markov to the data. Edge minimality in CPDAGs is discussed by Raskutti and Uhler, [13], Lam et al., [9]; here, we extend this to causally insufficient settings.
@@
For any two distinct nodes $A$ and $B$ in a maximal ancestral graph (MAG) $G$, the D-SEP set of $A$ relative to $B$, denoted $\text{D-SEP}(A,B)$, is the set of nodes $V\neq A$ such that there exists a collider path between $A$ and $V$ in which every node is an ancestor of either $A$ or $B$ in $G$. Intuitively, $\text{D-SEP}(A,B)$ contains the variables that may need to be conditioned on to separate $A$ and $B$ when latent confounding is present. In the oracle setting, if $A$ and $B$ are not adjacent in the true MAG, then there exists a subset of $\text{D-SEP}(A,B)$ that d-separates them [14].
@@
This set may include nonadjacent nodes and is used by FCI and GFCI to search for separating sets in the presence of latent structure [14].
@@
3 Simple Variants of GFCI: BOSS-FCI, GRaSP-FCI, and LV-Lite
The Greedy FCI (GFCI) algorithm [10] is a hybrid causal discovery method combining score-based search with conditional independence (CI) testing. It first estimates a Completed Partially Directed Acyclic Graph (CPDAG) using FGES, then refines it into a Partial Ancestral Graph (PAG) through FCI-style edge removal and orientation rules. This design reduces the search space while remaining robust to latent variables and selection bias.
@@ 3 Simple Variants of GFCI: BOSS-FCI, GRaSP-FCI, and LV-Lite
BOSS-FCI replaces FGES with the Best Order Score Search (BOSS) algorithm [2]. BOSS performs permutation-based score search, offering strong accuracy and scalability—especially for dense graphs. After estimating a CPDAG, the algorithm applies the standard GFCI refinement: adjacency-based edge removal, collider identification, and the FCI orientation rules.
@@
GRaSP-FCI substitutes the GRaSP algorithm [9] for FGES in the same framework. GRaSP employs a greedy adjacency search with penalty regularization, emphasizing sparsity and offering a complementary trade-off between accuracy and efficiency. Once the CPDAG is obtained, the usual GFCI refinement steps are applied.
LV-Lite.
LV-Lite [^3] LV-Lite has been referred to as BOSS-POD (BOSS PAG of DAG) elsewhere, but the name “LV-Lite” has stuck, so we defer. is a deliberately simple heuristic. Like BOSS-FCI, it begins with a CPDAG estimated by BOSS, but instead of performing any CI-based refinement, it directly outputs the PAG equivalence class of the resulting DAG. The method therefore captures latent-aware patterns implied by the BOSS structure but cannot remove edges to expose bidirected relationships. Although formally incomplete, LV-Lite is extremely fast and often achieves empirical accuracy comparable to its more principled counterparts.
@@
Appendix A provides pseudocode for the general *-FCI template, where the initial CPDAG procedure $\mathcal{M}$ may be FGES (standard GFCI), BOSS (BOSS-FCI), or GRaSP (GRaSP-FCI). LV-Lite can be viewed as a degenerate instance of this template, omitting edge removal and collider orientation and returning the PAG of the BOSS DAG directly.
@@ 4 The FCIT Algorithm
We now instantiate this FCI-Backward template as the FCIT (Fast Causal Inference Targeted-Testing) algorithm, designed to mitigate the repeated-testing problem in FCI-style searches, including other hybrid score-and-test methods. Following GFCI [10], FCIT begins with a score-based procedure that outputs a well-formed CPDAG—e.g., FGES, BOSS, GRaSP, or for small models SP. By default, we use BOSS for its scalability and accuracy.
As noted by Ogarrio et al., [10], score-based CPDAGs can include extra adjacencies or incorrect orientations when latent variables are present, since a CPDAG cannot represent bidirected edges. For instance, the MAG $X\to Y\leftrightarrow Z\leftarrow W$ may yield $X\to Y\to Z\leftarrow W$ (with an extra $X\to Z$) or $X\to Y\leftarrow Z\leftarrow W$ (with an extra $W\to Y$). GFCI corrects these by verifying independence ($X{\;\perp\!\!\!\perp\;}Z$ or $W{\;\perp\!\!\!\perp\;}Y$) and removing the spurious edge, restoring the correct collider structure. FCIT follows the same principle but uses targeted refinement—fewer, higher-value CI tests—while guaranteeing a sound PAG.
Algorithm 2 summarizes the FCIT procedure (see notation in Table 1). The initial CPDAG is converted to a PAG by replacing each edge with $\circ\!$-$\circ$ and transferring the unshielded colliders identified in the CPDAG. This is justified because the final PAG must be a subgraph of the CPDAG’s adjacencies, and collider orientations are the only ones that carry over reliably [10].

### v3

@@ Abstract
We present a family of score-guided mixed-strategy causal search algorithms that extend this framework. First, we introduce BOSS-FCI and GRaSP-FCI, variants of GFCI (Greedy Fast Causal Inference) that substitute BOSS (Best Order Score Search) or GRaSP (Greedy Relaxations of Sparsest Permutation) for FGES (Fast Greedy Equivalence Search), preserving correctness while trading off scalability and conservativeness. Second, we develop FCI Targeted-Testing (FCIT), a novel hybrid method that replaces exhaustive testing with targeted, score-informed tests guided by BOSS. FCIT guarantees well-formed PAGs and achieves higher precision and efficiency across sample sizes. Finally, we propose a lightweight heuristic, LV-Dumb (Latent Variable “Dumb”), which returns the PAG of the BOSS DAG (Directed Acyclic Graph). Though not strictly sound for latent confounding, LV-Dumb often matches FCIT’s accuracy while running substantially faster.
Simulations and real-data analyses show that BOSS-FCI and GRaSP-FCI provide robust baselines, FCIT yields the best balance of precision and reliability, and LV-Dumb offers a fast, near-equivalent alternative. Together, these methods demonstrate that targeted and score-guided strategies can dramatically improve the efficiency and correctness of latent-variable causal discovery.
@@ 1 Introduction
Learning causal structure from observational data is central in fields such as epidemiology, economics, and biology, where randomized experiments are often infeasible. In many real-world settings, latent variables or selection bias complicate this task, undermining the reliability of traditional discovery methods. The Fast Causal Inference (FCI) algorithm [spirtes2001causation] provides a principled framework for such cases by applying conditional independence (CI) tests to subsets of adjacent variables (or Possible-D-SEP sets). While correct in theory, this strategy is fragile in practice: testing many subsets amplifies stochastic variation in $p$-values, leading to spurious independence claims, missing or extra edges, and unstable orientations. We refer to this as the repeated testing problem.
@@
Beyond sampling variability, another challenge is almost unfaithfulness [zhang2008detection, spirtes2014uniformly]. Even when the Causal Markov and Faithfulness assumptions hold exactly, small parameter perturbations can produce distributions that are nearly unfaithful—where partial correlations or other dependence measures are close to zero but not exactly zero. In finite samples, such “almost unfaithful” parameterizations lead to unstable independence decisions: small statistical fluctuations can flip dependence judgments, causing orientation errors or spurious adjacencies. As shown by zhang2008detection, this phenomenon can make reliable causal inference difficult even for consistent algorithms, motivating procedures that are robust to weak dependencies or employ stronger assumptions such as $k$-triangle faithfulness [spirtes2014uniformly]. We therefore treat almost unfaithfulness as a complementary concern to the repeated-testing problem—structural rather than statistical in origin, but equally capable of producing unstable PAGs in practice.
@@
Several extensions have sought to mitigate such issues. Order-independent FCI variants [colombo2014order] stabilize orientations, and score-based methods [^1] Scores in Tetrad currently cover linear Gaussian and linear non-Gaussian models (using BIC-type criteria), multinomial models, mixed multinomial/linear Gaussian models, and, more recently, mixed multinomial/nonlinear post-nonlinear models via truncated basis-function scores [ramsey2025scalablecausaldiscoveryrecursive]. extend causal search to latent-variable models through likelihood-based criteria [triantafillou2016score, bhattacharya2021differentiable]. Hybrid approaches such as Greedy FCI (GFCI; [ogarrio2016hybrid]) combine score-based search with CI testing, but still rely on adjacency-based phases—i.e., during the adjacency-removal stage, testing all subsets of nodes adjacent to a pair—and thus remain vulnerable to repeated testing and almost unfaithfulness. Constraint-based methods, by contrast, have the advantage of being inherently nonparametric and making no distributional assumptions except for those of the selected conditional independence test, which motivates the option of keeping such testing in the algorithm.
@@
Recent advances in score-based search motivate new hybrids that combine the distributional flexibility of constraint-based methods with the efficiency of likelihood-based optimization. Best Order Score Search (BOSS; [andrews2019learning]) provides a scalable, permutation-based alternative to FGES, while GRaSP [lam2022greedy] offers a related variant. Substituting these into the GFCI framework yields BOSS-FCI and GRaSP-FCI—hybrid algorithms that preserve the nonparametric generality of constraint-based search while leveraging score-based initialization to improve stability and scalability. The original GFCI algorithm [ogarrio2016hybrid] is sound and complete in the population limit under the Causal Markov and Faithfulness assumptions, returning the correct Partial Ancestral Graph when provided with perfect independence information. However, in finite samples, GFCI and its variants still depend on adjacency-based conditioning and thus remain susceptible to repeated testing and almost unfaithfulness.
@@
Building on these baselines, we propose FCI Targeted-Testing (FCIT), a mixed-strategy algorithm that replaces exhaustive adjacency-based testing with a path-based procedure. Like GFCI, FCIT begins with a score-based completed partially directed acyclic graph (CPDAG)—a representation of the Markov equivalence class of DAGs that share the same adjacencies and unshielded colliders. Score-based algorithms such as BOSS and GRaSP are score-consistent: in the large-sample limit, they recover the correct CPDAG under the Causal Markov and Faithfulness assumptions for the model class (e.g., linear Gaussian, multinomial, or post-nonlinear) used to define the score. FCIT then applies a recursive routine, block_paths_recursively, to identify conditioning sets only when necessary to block active paths between variables, thereby avoiding the over-conditioning that arises in FCI and GFCI from testing all subsets of adjacent variables. Compared with these algorithms, FCIT performs far fewer conditional independence tests, which reduces sampling error and increases the reliability of the overall learned structure. Together with a well-formedness check that enforces the rules of a Partial Ancestral Graph (PAG)—a mixed graph representing the Markov equivalence class of maximal ancestral graphs consistent with possible latent variables and selection bias—this targeted strategy guarantees that every intermediate graph remains structurally consistent, meaning it obeys the ancestral, acyclicity, and maximality constraints required of a valid PAG. The result is greater efficiency, more stable orientations, and structural correctness not ensured by prior FCI variants [zhang2008completeness, hu2024fast].
@@
Finally, we introduce a deliberately simple heuristic, LV-Dumb (Latent Variables “Dumb”), which outputs the PAG implied by the BOSS DAG without performing any latent-specific conditional independence testing. Though it cannot orient bidirected edges, LV-Dumb scales exceptionally well and, in extensive simulation studies, often matches or exceeds the empirical accuracy of more complex algorithms when compared to the true PAG.
@@
Together, these four algorithms—BOSS-FCI, GRaSP-FCI, FCIT, and LV-Dumb—form a continuum. The first two establish sound baselines by embedding advanced score-based methods within the GFCI framework. FCIT introduces a principled path-based refinement that improves efficiency and reliability, while LV-Dumb represents a heuristic extreme: formally incomplete but practically effective. Our empirical evaluation, spanning both simulation and real data, demonstrates the complementary strengths of these approaches. BOSS-FCI and GRaSP-FCI retain theoretical correctness but are constrained by repeated testing; FCIT balances correctness with improved precision and scalability; and LV-Dumb sacrifices correctness guarantees yet achieves strong accuracy with minimal computational cost in simulation.
@@
The remainder of this paper is organized as follows. Section 2 reviews graphical preliminaries. Sections 34 present the four algorithms. Section 5 introduces optimization details, Section 6 outlines theoretical guarantees, and Section 8 reports empirical results. All algorithms are implemented in the open-source Tetrad suite [ramsey2018tetrad], with reproducibility materials available on GitHub.
@@
  • A heuristic, LV-Dumb, that outputs the PAG of the BOSS DAG, achieving high empirical accuracy at minimal cost.
@@ 2 Graphical Preliminaries
In this section, we define key terms related to graphical causal models used throughout the paper. We focus on models based on causal directed acyclic graphs (DAGs), which consist of directed edges $X\rightarrow Y$. Such an edge indicates that variable $X$ causes variable $Y$, meaning that interventions on $X$ counterfactually alter $Y$ [spirtes2001causation].
@@
A MAG or PAG is maximal (following Zhang [zhang2008completeness]). if, for every non-adjacent pair $X,Y$, there exists a conditioning set S such that $X$ and $Y$ are m-separated given S. Equivalently, there are no inducing paths between non-adjacent nodes.
@@
A set of variables $\textbf{(}V)$ is causally sufficient if it includes all common causes of observed variables in $\textbf{(}V)$; otherwise, it is causally insufficient. A model is edge-minimal of a certain type (e.g., DAG, CPDAG, PAG) if no model of that type has fewer edges while remaining Markov to the data. Edge minimality in CPDAGs is discussed by raskutti2018learning, lam2022greedy; here, we extend this to causally insufficient settings.
@@
For any two distinct nodes $A$ and $B$ in a maximal ancestral graph (MAG) $G$, the D-SEP set of $A$ relative to $B$, denoted $\text{D-SEP}(A,B)$, is the set of nodes $V\neq A$ such that there exists a collider path between $A$ and $V$ in which every node is an ancestor of either $A$ or $B$ in $G$. Intuitively, $\text{D-SEP}(A,B)$ contains the variables that may need to be conditioned on to separate $A$ and $B$ when latent confounding is present. In the oracle setting, if $A$ and $B$ are not adjacent in the true MAG, then there exists a subset of $\text{D-SEP}(A,B)$ that d-separates them [spirtes2001causation].
@@
This set may include nonadjacent nodes and is used by FCI and GFCI to search for separating sets in the presence of latent structure [spirtes2001causation].
@@
3 Simple Variants of GFCI: BOSS-FCI, GRaSP-FCI, and LV-Dumb
The Greedy FCI (GFCI) algorithm [ogarrio2016hybrid] is a hybrid causal discovery method combining score-based search with conditional independence (CI) testing. It first estimates a Completed Partially Directed Acyclic Graph (CPDAG) using FGES, then refines it into a Partial Ancestral Graph (PAG) through FCI-style edge removal and orientation rules. This design reduces the search space while remaining robust to latent variables and selection bias.
@@ 3 Simple Variants of GFCI: BOSS-FCI, GRaSP-FCI, and LV-Dumb
BOSS-FCI replaces FGES with the Best Order Score Search (BOSS) algorithm [andrews2019learning]. BOSS performs permutation-based score search, offering strong accuracy and scalability—especially for dense graphs. After estimating a CPDAG, the algorithm applies the standard GFCI refinement: adjacency-based edge removal, collider identification, and the FCI orientation rules.
@@
GRaSP-FCI substitutes the GRaSP algorithm [lam2022greedy] for FGES in the same framework. GRaSP employs a greedy adjacency search with penalty regularization, emphasizing sparsity and offering a complementary trade-off between accuracy and efficiency. Once the CPDAG is obtained, the usual GFCI refinement steps are applied.
LV-Dumb.
LV-Dumb [^3] LV-Dumb has been referred to as BOSS-POD (BOSS PAG of DAG) elsewhere, but the name “LV-Dumb” has stuck, so we defer. is a deliberately simple heuristic. Like BOSS-FCI, it begins with a CPDAG estimated by BOSS, but instead of performing any CI-based refinement, it directly outputs the PAG equivalence class of the resulting DAG. The method therefore captures latent-aware patterns implied by the BOSS structure but cannot remove edges to expose bidirected relationships. Although formally incomplete, LV-Dumb is extremely fast and often achieves empirical accuracy comparable to its more principled counterparts.
@@
Appendix A provides pseudocode for the general *-FCI template, where the initial CPDAG procedure $\mathcal{M}$ may be FGES (standard GFCI), BOSS (BOSS-FCI), or GRaSP (GRaSP-FCI). LV-Dumb can be viewed as a degenerate instance of this template, omitting edge removal and collider orientation and returning the PAG of the BOSS DAG directly.
@@ 4 The FCIT Algorithm
We now instantiate this FCI-Backward template as the FCIT (Fast Causal Inference Targeted-Testing) algorithm, designed to mitigate the repeated-testing problem in FCI-style searches, including other hybrid score-and-test methods. Following GFCI [ogarrio2016hybrid], FCIT begins with a score-based procedure that outputs a well-formed CPDAG—e.g., FGES, BOSS, GRaSP, or for small models SP. By default, we use BOSS for its scalability and accuracy.
As noted by ogarrio2016hybrid, score-based CPDAGs can include extra adjacencies or incorrect orientations when latent variables are present, since a CPDAG cannot represent bidirected edges. For instance, the MAG $X\to Y\leftrightarrow Z\leftarrow W$ may yield $X\to Y\to Z\leftarrow W$ (with an extra $X\to Z$) or $X\to Y\leftarrow Z\leftarrow W$ (with an extra $W\to Y$). GFCI corrects these by verifying independence ($X{\;\perp\!\!\!\perp\;}Z$ or $W{\;\perp\!\!\!\perp\;}Y$) and removing the spurious edge, restoring the correct collider structure. FCIT follows the same principle but uses targeted refinement—fewer, higher-value CI tests—while guaranteeing a sound PAG.
Algorithm 2 summarizes the FCIT procedure (see notation in Table 1). The initial CPDAG is converted to a PAG by replacing each edge with $\circ\!$-$\circ$ and transferring the unshielded colliders identified in the CPDAG. This is justified because the final PAG must be a subgraph of the CPDAG’s adjacencies, and collider orientations are the only ones that carry over reliably [ogarrio2016hybrid].